write the relation between phase and group velocity
1
Expert's answer
2020-09-15T11:29:54-0400

Phase velocity(wave velocity) is the velocity of monochromatic wave.Velocity propagated when a single wave of definite wave length travel in a medium.

Given by;

"V\scriptscriptstyle p" ="\dfrac{\omega}{k}" where k is a constant.

Group velocity: velocity of a number of wave of different wavelength moving with a different velocity in medium.

Given by:

"V\scriptscriptstyle g" ="\dfrac{d\scriptscriptstyle \omega}{d\scriptscriptstyle k}"

Where wave frequency is given by;

"\omega" =k"V\scriptscriptstyle p"

Differentiating wave frequency with respect to k;

"V\scriptscriptstyle g" ="\dfrac{d }{d\scriptscriptstyle k}(" "KV\scriptscriptstyle p" )

"V\scriptscriptstyle g" = "V\scriptscriptstyle p" +"k\dfrac{dV\scriptscriptstyle p}{d\scriptscriptstyle k}"

Taking k ="\dfrac{2\pi }{\lambda}"

"\because" "\dfrac{d\scriptscriptstyle k}{d\scriptscriptstyle\lambda}" ="-\dfrac{2\pi }{\lambda^2}"

"d\scriptscriptstyle k" ="-k\dfrac{d\lambda }{\lambda}"

"V\scriptscriptstyle g" ="V\scriptscriptstyle p" "-\lambda\dfrac{dV\scriptscriptstyle p }{d\lambda}"